e TOnSWing array: intl] a- 1,3,5,7,9,11 (2x2 marks) What is the value stored in the variable total when the followings loops complete? Assume there are ho errors in the code. Assume the loops are independent (L.e., part b is not affected by part a). a) int total-0 f total-total + alil: b) int total = 0; for (inti-l; i<5; i+-2) total total + afi]i
Answer #1
a)
total stores value of a[0] + a[1] + a[2] + a[3] + a[4] + a[5] = 1 + 3 + 5 + 7 + 9 + 11 = 36
so, total stores a value of 36.

b)
total stores value of a[1] + a[3] = 3 + 7 = 10
so, total stores a value of 10.